# An Elliott intertwining approach to classifying actions of $\mathrm{C}^{*}$-tensor categories

### 29 Jan 2024 v2

@@ 4.1. Topology induced on cocycle representations
Note that $F(1_{\mathcal{C}})=A$ and for any $a\in A$ and $b\in B$ $\mathbbm{v}_{1_{\mathcal{C}}}(a\boxtimes_{\phi}b)=\phi(a)b$ and $\mathbbm{w}_{1_{\mathcal{C}}}(a\boxtimes_{\psi}b)=\psi(a)b$. Therefore, the family of pseudometrics $d_{\mathcal{F}}$ bounds the pointwise distance between $\phi$ and $\psi$ in the strict topology.
@@
Suppose that for any $X\in\mathrm{Irr}(\mathcal{C})$, $h_{\lambda}^{X}$ converges pointwise to $h^{X}$ in the norm induced by the right inner product. Let $\zeta_{\mu}$ be an approximate unit of $A$ and recall from (3.10) that for any $X\in\mathcal{C}$, $x\in F(X)$, and $b\in B$
@@
In [40, Definition 2.5], Szabó defines a topology on cocycle morphisms, which using the conventions of this paper, is generated by the family of pseudometrics
@@
The convergence with respect to the family of pseudometrics in (4.14) implies convergence with respect to the family of pseudometrics in (4.12). If $\phi$ and $\psi$ are non-degenerate, then $\phi(A)B$ and $\psi(A)B$ are dense in $B$. Moreover, the case $g=1$ in (4.12) recovers the pointwise difference of the morphisms. Thus, convergence with respect to the family of pseudometrics in (4.12) implies convergence with respect to the family of pseudometrics in (4.14). However, these topologies are different outside the non-degenerate setting, with the topology induced by the family of pseudometrics in (4.12) being coarser than the topology induced by the family of pseudometrics in (4.14). Finally, we would like to point out that the topology induced by (4.13) coincides with that used in [31].
@@
Proof.
Let $(\phi_{\lambda},h_{\lambda}):(A,F,J)\to(B,G,I)$ and $(\psi_{\lambda},l_{\lambda}):(B,G,I)\to(C,H,K)$ be two convergent nets with limits $(\phi,h)$ and $(\psi,l)$ respectively. First, note that for any $X\in\mathrm{Irr}(\mathcal{C})$, $l^{X}\circ h^{X}=\lim\limits_{\lambda\to\infty}l_{\lambda}^{X}\circ h_{\lambda}^{X}$ in the point-norm topology. Then, the case $X=1_{\mathcal{C}}$ yields that $\psi\circ\phi=\lim\limits_{\lambda\to\infty}\psi_{\lambda}\circ\phi_{\lambda}$ holds in the point-norm topology. Using the composition formula in Lemma 3.15 and Lemma 4.9, the conclusion follows. ∎
4.2. Approximate unitary equivalence
@@
Suppose $\mathcal{C}$ is a semisimple $\mathrm{C}^{*}$-tensor category with countably many isomorphism classes of simple objects acting on a $\mathrm{C}^{*}$-algebra $B$. We denote this action by the triple $(B,G,I)$. For any $u$ unitary in $\mathcal{M}(B)$, we consider $\mathrm{Ad}(u):B\to B$ to be the $*{}^{*}$-homomorphism given by $b\mapsto ubu^{*}$. Then $\mathrm{Ad}(u)$ induces a Hilbert $B$-bimodule ${}_{\mathrm{Ad}(u)}B$. The map $T_{u}:{}_{\mathrm{Ad}(u)}B\to B$ given by $b\mapsto u^{*}b$ is a bimodule isomorphism. But for any $X\in\mathcal{C}$, $G(X)\boxtimes B\cong B\boxtimes G(X)$, so there exists a unitary isomorphism $(\mathbbm{v}_{u})_{X}:G(X)\boxtimes{}_{\mathrm{Ad}(u)}B\to{}_{\mathrm{Ad}(u)}B\boxtimes G(X)$. It follows that $(\mathrm{Ad}(u),\mathbbm{v}_{u}):(B,G,I)\to(B,G,I)$ is a cocycle morphism. We denote by $h_{u}=\{h_{u}^{X}:G(X)\to G(X)\}_{X\in\mathcal{C}}$ the collection of linear maps corresponding to the cocycle morphism induced by $\mathrm{Ad}(u)$.
@@
For any $X\in\mathcal{C}$, consider the bimodule maps $L_{X}:G(X)\to B\boxtimes G(X)$ and $R_{X}:G(X)\to G(X)\boxtimes B$ given by $L_{X}(x)=\lim\limits_{\mu}\eta_{\mu}\boxtimes x$ and $R_{X}(x)=\lim\limits_{\mu}x\boxtimes\eta_{\mu}$ for some $\eta_{\mu}$ quasicentral approximate unit of $B$ with respect to $\mathcal{M}(B)$.
@@
$\iota:G(X)\to G(X)\boxtimes{}_{\mathrm{Ad}(u)}B\to G(X)\boxtimes B$
given by $x\mapsto\lim\limits_{\mu}x\boxtimes\eta_{\mu}\mapsto\lim\limits_{\mu}x\boxtimes u^{*}\eta_{\mu}$. Since $\eta_{\mu}$ is quasicentral,
@@
[…]
@@
    Let $(\phi,\mathbbm{v}):(A,F,J)\to(B,G,I)$ be a cocycle representation and $u\in\mathcal{M}(B)$ a unitary. We say that $(\psi,\mathbbm{w})$ is an approximate unitary conjugate of $(\phi,\mathbbm{v})$ if there exists a net of unitaries $u_{\lambda}\in\mathcal{U}(\mathcal{M}(B))$ such that
@@
[…]
@@
  3. (iii)
    We say that $(\phi,\mathbbm{v})$ and $(\psi,\mathbbm{w})$ are approximately unitarily equivalent if $(\psi,\mathbbm{w})\precapprox_{u}(\phi,\mathbbm{v})$ and $(\phi,\mathbbm{v})\precapprox_{u}(\psi,\mathbbm{w})$. This will be denoted by $(\phi,\mathbbm{v})\approx_{u}(\psi,\mathbbm{w})$.
In general, we will be interested in these notions when $(\phi,\mathbbm{v})$ and $(\psi,\mathbbm{w})$ are cocycle morphisms, so let us record the following lemma.
@@
If $(\phi,\mathbbm{v}),(\psi,\mathbbm{w}):(A,F,J)\to(B,G,I)$ are cocycle morphisms with associated families of linear maps $\{h^{X}\}$ and $\{l^{X}\}$ respectively, then $(\psi,\mathbbm{w})$ is an approximate unitary conjugate of $(\phi,\mathbbm{v})$ if and only if there exists a net of unitaries $u_{\lambda}\in\mathcal{U}(\mathcal{M}(B))$ such that
[…]
@@
Proof.
By definition, $(\psi,\mathbbm{w})$ is an approximate unitary conjugate of $(\phi,\mathbbm{v})$ if and only if there exists a net of unitaries $u_{\lambda}$ in $\mathcal{M}(B)$ inducing a net of cocycle morphisms $(\mathrm{Ad}(u_{\lambda}),\mathbbm{v}_{u_{\lambda}})$ such that $(\mathrm{Ad}(u_{\lambda}),\mathbbm{v}_{u_{\lambda}})\circ(\phi,\mathbbm{v})$ converges to $(\psi,\mathbbm{w})$. Suppose the cocycle morphisms $(\mathrm{Ad}(u_{\lambda}),\mathbbm{v}_{u_{\lambda}})$ have associated families of linear maps $\{h_{u_{\lambda}}^{X}\}$. By Lemma 4.9, $(\mathrm{Ad}(u_{\lambda}),\mathbbm{v}_{u_{\lambda}})\circ(\phi,\mathbbm{v})$ converges to $(\psi,\mathbbm{w})$ if and only if $h_{u_{\lambda}}^{X}\circ h^{X}$ converges pointwise to $l^{X}$ for any $X\in\mathrm{Irr}(\mathcal{C})$. Moreover, by Lemma 4.12,
$h_{u_{\lambda}}^{X}(h^{X}(x))=u_{\lambda}\rhd h^{X}(x)\lhd u_{\lambda}^{*}$
for all $X\in\mathcal{C}$ and all $x\in F(X)$, so the conclusion follows. ∎
Remark 4.15.
There is a key difference to Szabó’s notion of approximate unitary conjugacy in [40, Definition 2.8]. The notion of approximate unitary conjugacy in Definition 4.13 is symmetric, even when restricted to possibly degenerate morphisms. In Lemma 4.16, we show that approximate unitary conjugacy implies approximate unitary equivalence. This is a consequence of the difference in topologies (see Example 4.10).
Lemma 4.16.
Let $(\phi,h),(\psi,l):(A,F,J)\to(B,G,I)$ be cocycle morphisms. If $(\psi,l)$ is an approximate unitary conjugate of $(\phi,h)$, then $(\phi,h)$ and $(\psi,l)$ are approximately unitarily equivalent.
Proof.
Suppose there exist unitaries $u_{\lambda}\in\mathcal{U}(\mathcal{M}(B))$ such that
[…]
for any $X\in\mathrm{Irr}(\mathcal{C})$, and any $x\in F(X)$. We show that
[…]
for any $X\in\mathrm{Irr}(\mathcal{C})$, and any $x\in F(X)$.
Note that for any $X\in\mathcal{C}$ and any $x\in F(X)$, we have that
$h_{u_{\lambda}^{*}}^{X}(h_{u_{\lambda}}^{X}(h^{X}(x)))=u_{\lambda}^{*}\rhd h_{u_{\lambda}}^{X}(h^{X}(x))\lhd u_{\lambda}=1_{\mathcal{M}(B)}\rhd h^{X}(x)\lhd 1_{\mathcal{M}(B)}.$
But $G(X)$ is non-degenerate, so $1_{\mathcal{M}(B)}\rhd h^{X}(x)\lhd 1_{\mathcal{M}(B)}=h^{X}(x)$ by Lemma 1.11. Therefore, for any $X\in\mathcal{C}$ and any $x\in F(X)$,
$\|h_{u_{\lambda}^{*}}^{X}\circ l^{X}(x)-h^{X}(x)\|=\|h_{u_{\lambda}^{*}}^{X}(l^{X}(x)-h_{u_{\lambda}}^{X}\circ h^{X}(x))\|.$
Take $y_{\lambda}=l^{X}(x)-h_{u_{\lambda}}^{X}\circ h^{X}(x)$ and note that
$\langle h_{u_{\lambda}^{*}}^{X}(y_{\lambda}),h_{u_{\lambda}^{*}}^{X}(y_{\lambda})\rangle_{B}=\mathrm{Ad}(u_{\lambda}^{*})(\langle y_{\lambda},y_{\lambda}\rangle_{B}).$
[…]
@@
[…]
@@
$\|\mathbbm{w}_{X}(x\boxtimes_{\psi}b)-(\mathbbm{v}_{u_{t}}*\mathbbm{v})_{X}(x\boxtimes_{\phi}b)\|\overset{t\to\infty}{\longrightarrow}0$
for all $a\in A$, $b\in B$, $X\in\mathrm{Irr}(\mathcal{C})$, and any $x\in F(X)$. This will be denoted by $(\phi,\mathbbm{v})\cong_{u}(\psi,\mathbbm{w})$. [^15] Following Definition 4.13, we could have defined asymptotic unitary conjugacy. However, this implies asymptotic unitary equivalence as in Lemma 4.16.
The same argument as in Lemma 4.14 gives the following equivalent characterisation for asymptotic unitary equivalence of cocycle morphisms.
@@
Let $(\phi_{1},h_{1}):(A,F,J)\to(B,G,I),(\psi_{1},l_{1}):(B,G,I)\to(C,H,K)$ be two cocycle morphisms. Suppose that $(\phi_{1},h_{1})$ is unitarily equivalent to $(\phi_{2},h_{2})$ and $(\psi_{1},l_{1})$ is unitarily equivalent to $(\psi_{2},l_{2})$. If $\psi_{1}$ is extendible, then $(\psi_{1}\circ\phi_{1},l_{1}\circ h_{1})$ is unitarily equaivalent to $(\psi_{2}\circ\phi_{2},l_{2}\circ h_{2})$. Indeed, let $u\in\mathcal{U}(\mathcal{M}(B))$ and $v\in\mathcal{U}(\mathcal{M}(C))$ be such that $u\rhd h_{1}^{X}(x)\lhd u^{*}=h_{2}^{X}(x)$ and $v\rhd l_{1}^{X}(y)\lhd v^{*}=l_{2}^{X}(y)$ for any $X\in\mathcal{C}$, $x\in F(X)$, and $y\in G(X)$. Then, $v\psi_{1}^{\dagger}(u)\rhd l_{1}^{X}(h_{1}^{X}(x))\lhd(v\psi_{1}^{\dagger}(u))^{*}=l_{2}^{X}(h_{2}^{X}(x))$. In general, the vertical composition in the $2$-category of correspondences is not well defined outside of the extendible setting.

### 08 Apr 2026 v3

@@ 5. Approximate unitary equivalence
Note that $F(1_{\mathcal{C}})=A$ and for any $a\in A$ and $b\in B$ $\mathbbm{v}_{1_{\mathcal{C}}}(a\boxtimes_{\phi}b)=\phi(a)b$ and $\mathbbm{w}_{1_{\mathcal{C}}}(a\boxtimes_{\psi}b)=\psi(a)b$.
@@
Suppose that for any $X\in\mathrm{Irr}(\mathcal{C})$, $h_{\lambda}^{X}$ converges pointwise to $h^{X}$ in the norm induced by the right inner product. Let $\zeta_{\mu}$ be an approximate unit of AA and recall from (3.5) that for any $X\in\mathcal{C}$, $x\in F(X)$, and $b\in B$
@@
In [41, Definition 2.5], Szabó defines a topology on the space of cocycle morphisms which is generated by the family of pseudometrics
@@
The convergence with respect to the family of pseudometrics in (5.6) implies convergence with respect to the family of pseudometrics in (5.4). If $\phi$ and $\psi$ are non-degenerate, then $\phi(A)B$ and $\psi(A)B$ are dense in $B$. Moreover, the case $g=1$ in (5.4) recovers the pointwise difference of the morphisms. Thus, convergence with respect to the family of pseudometrics in (5.4) implies convergence with respect to the family of pseudometrics in (5.6). However, these topologies are different outside the non-degenerate setting, with the topology induced by the family of pseudometrics in (5.4) being coarser than the topology induced by the family of pseudometrics in (5.6). Finally, we would like to point out that the topology induced by (5.5) coincides with that used in [33, Section 5].
@@
Suppose $\mathcal{C}$ is a semisimple $\mathrm{C}^{*}$-tensor category with countably many isomorphism classes of simple objects acting on a $\mathrm{C}^{*}$-algebra $B$. We denote this action by the triple $(B,G,I)$. For any $u$ unitary in $\mathcal{M}(B)$, we consider $\mathrm{Ad}(u):B\to B$ to be the -homomorphism given by $b\mapsto ubu^{*}$. Then $\mathrm{Ad}(u)$ induces a Hilbert $B$-bimodule $_{\mathrm{Ad}(u)}B$. The map $T_{u}:{}_{\mathrm{Ad}(u)}B\to B$ given by $b\mapsto u^{*}b$ is a bimodule isomorphism. But for any $X\in\mathcal{C}$, $G(X)\boxtimes B\cong B\boxtimes G(X)$, so there exists a unitary isomorphism $(\mathbbm{v}_{u})_{X}:G(X)\boxtimes{}_{\mathrm{Ad}(u)}B\to{}_{\mathrm{Ad}(u)}B\boxtimes G(X)$. It follows that $(\mathrm{Ad}(u),\mathbbm{v}_{u}):(B,G,I)\to(B,G,I)$ is a cocycle morphism.. We denote by $h_{u}=\{h_{u}^{X}:G(X)\to G(X)\}_{X\in\mathcal{C}}$ the collection of linear maps corresponding to the cocycle morphism induced by $\mathrm{Ad}(u)$.
@@
For any $X\in\mathcal{C}$, consider the bimodule maps $L_{X}:G(X)\to B\boxtimes G(X)$ and $R_{X}:G(X)\to G(X)\boxtimes B$ given by $L_{X}(x)=\lim\limits_{\mu}\eta_{\mu}\boxtimes x$ and $R_{X}(x)=\lim\limits_{\mu}x\boxtimes\eta_{\mu}$ for some $\eta_{\mu}$ quasicentral approximate unit of $B$ with respect to $\mathcal{M}(B)$ (i.e. $\eta_{\mu}x-x\eta_{\mu}\to 0$ for all $x\in\mathcal{M}(B)$).
@@
$\iota:G(X)\to G(X)\boxtimes B$
given by $x\mapsto\lim\limits_{\mu}x\boxtimes u^{*}\eta_{\mu}$. Since $\eta_{\mu}$ is quasicentral,
@@
Let $\mathcal{C}\overset{F}{\curvearrowright}A$ and $\mathcal{C}\overset{G}{\curvearrowright}B$ be actions of $\mathcal{C}$ on $\mathrm{C}^{*}$-algebras AA and $B$ and let $(\phi,\mathbbm{v}),(\psi,\mathbbm{w}):(A,F,J)\to(B,G,I)$ be cocycle representations.
@@
    We say that the pairs $(\phi,\mathbbm{v})$ and $(\psi,\mathbbm{w})$ are unitarily equivalent if there exists a unitary $u\in\mathcal{U}(\mathcal{M}(B))$ such that
@@
    We say that $(\psi,\mathbbm{w})$ is an approximate unitary conjugate of $(\phi,\mathbbm{v})$ if there exists a net of unitaries $u_{\lambda}\in\mathcal{U}(\mathcal{M}(B))$ such that
@@
    $\|\mathbbm{w}_{X}(x\boxtimes_{\psi}b)-(\mathbbm{v}_{u_{\lambda}}*\mathbbm{v})_{X}(x\boxtimes_{\mathrm{Ad}(u_{\lambda})\phi}b)\|\overset{\lambda}{\longrightarrow}0$
@@
The analogous notion of approximate unitary conjugacy in [41, Definition 2.8] is only a subequivalence relation in general. Indeed, symmetry fails when the cocycle representations considered are not extendible ([41, Remark 2.9]). However, the topology that we consider does indeed make it an equivalence relation.
@@
Let $(\phi,\mathbbm{v}),(\psi,\mathbbm{w}):(A,F,J)\rightarrow(B,G,I)$ be cocycle representations between $\mathcal{C}$-C-algebras. Then if $(\psi,\mathbbm{w})\precapprox_{u}(\phi,\mathbbm{v})$, one also has that $(\phi,\mathbbm{v})\precapprox_{u}(\psi,\mathbbm{w})$.
Proof.
It can be computed from the definitions that for any unitary $u\in\mathcal{U}(\mathcal{M}(B))$, $X\in\mathcal{C}$ and $x\in F(X)$
$(\mathbbm{v}_{u}*\mathbbm{v})_{X}(x\boxtimes_{\mathrm{Ad}(u)\phi}b)=u\rhd\mathbbm{v}_{X}(x\boxtimes_{\phi}u^{*}b).$
In particular
$\|\mathbbm{w}_{X}(x\boxtimes_{\psi}b)-(\mathbbm{v}_{u_{\lambda}}*\mathbbm{v})_{X}(x\boxtimes_{\mathrm{Ad}(u_{\lambda})\phi}b)\|$
$=\lVert\mathbbm{w}_{X}(x\boxtimes_{\psi}b)-u_{\lambda}\rhd\mathbbm{v}_{X}(x\boxtimes_{\phi}u_{\lambda}^{*}b)\rVert$
$=\lVert u_{\lambda}^{*}\rhd\mathbbm{w}_{X}(x\boxtimes_{\psi}b)-\mathbbm{v}_{X}(x\boxtimes_{\phi}u_{\lambda}^{*}b)\rVert$
$=\lVert(\mathbbm{v}_{u_{\lambda}^{*}}*\mathbbm{w})_{X}(x\boxtimes_{\mathrm{Ad}(u_{\lambda}^{*})\psi}u_{\lambda}^{*}b)-\mathbbm{v}_{X}(x\boxtimes_{\phi}u_{\lambda}^{*}b)\rVert.$
Thus the result follows. ∎
In light of Lemma 5.9, we simply call two cocycle representations
$(\psi,\mathbbm{w}),(\phi,\mathbbm{v}):(A,F,J)\rightarrow(B,G,I)$
such that $(\psi,\mathbbm{w})\precapprox_{u}(\phi,\mathbbm{v})$ approximately unitarily equivalent and we denote it rather by $(\phi,\mathbbm{v})\approx_{u}(\psi,\mathbbm{w})$. In general, we will be interested in these notions when $(\phi,\mathbbm{v})$ and $(\psi,\mathbbm{w})$ are cocycle morphisms, so let us record the following immediate consequence of Lemma 5.7.
Lemma 5.10.
If $(\phi,\mathbbm{v}),(\psi,\mathbbm{w}):(A,F,J)\to(B,G,I)$ are cocycle morphisms with associated families of linear maps $\{h^{X}\}$ and $\{l^{X}\}$ respectively, then $(\psi,\mathbbm{w})$ is approximately unitarily equivalent to $(\phi,\mathbbm{v})$ if and only if there exists a net of unitaries $u_{\lambda}\in\mathcal{U}(\mathcal{M}(B))$ such that
$\|l^{X}(x)-u_{\lambda}\rhd h^{X}(x)\lhd u_{\lambda}^{*}\|\overset{\lambda}{\longrightarrow}0$
@@
Let $\mathcal{C}\overset{F}{\curvearrowright}A$ and $\mathcal{C}\overset{G}{\curvearrowright}B$ be actions of $\mathcal{C}$ on $\mathrm{C}^{*}$-algebras AA and $B$ respectively and let $(\phi,\mathbbm{v}),(\psi,\mathbbm{w}):(A,F,J)\to(B,G,I)$ be cocycle representations. We say that $(\psi,\mathbbm{w})$ is asymptotically unitarily equivalent to $(\phi,\mathbbm{v})$ if there exists a strictly continuous map $u:[0,\infty)\to\mathcal{U}(\mathcal{M}(B))$ such that
@@
$\|\mathbbm{w}_{X}(x\boxtimes_{\psi}b)-(\mathbbm{v}_{u_{t}}*\mathbbm{v})_{X}(x\boxtimes_{\mathrm{Ad}(u_{t})\phi}b)\|\overset{t\to\infty}{\longrightarrow}0$
for all $a\in A$, $b\in B$, $X\in\mathrm{Irr}(\mathcal{C})$, and any $x\in F(X)$. This will be denoted by $(\phi,\mathbbm{v})\cong_{u}(\psi,\mathbbm{w})$.
The same argument as in Lemma 5.9 shows that asymptotic unitary equivalence is an equivalence relation. Moreover, the same argument as in Lemma 5.10 gives the following equivalent characterisation for asymptotic unitary equivalence of cocycle morphisms.
@@
Let $(\phi_{1},h_{1}):(A,F,J)\to(B,G,I),(\psi_{1},l_{1}):(B,G,I)\to(C,H,K)$ be two cocycle morphisms. Suppose that $(\phi_{1},h_{1})$ is unitarily equivalent to $(\phi_{2},h_{2})$ and $(\psi_{1},l_{1})$ is unitarily equivalent to $(\psi_{2},l_{2})$. If $\psi_{1}$ is extendible, then $(\psi_{1}\circ\phi_{1},l_{1}\circ h_{1})$ is unitarily equaivalent to $(\psi_{2}\circ\phi_{2},l_{2}\circ h_{2})$. Indeed, let $u\in\mathcal{U}(\mathcal{M}(B))$ and $v\in\mathcal{U}(\mathcal{M}(C))$ be such that $u\rhd h_{1}^{X}(x)\lhd u^{*}=h_{2}^{X}(x)$ and $v\rhd l_{1}^{X}(y)\lhd v^{*}=l_{2}^{X}(y)$ for any $X\in\mathcal{C}$, $x\in F(X)$, and $y\in G(X)$. Then, $v\psi_{1}^{\dagger}(u)\rhd l_{1}^{X}(h_{1}^{X}(x))\lhd(v\psi_{1}^{\dagger}(u))^{*}=l_{2}^{X}(h_{2}^{X}(x))$. In general, the vertical composition in the 22-category of correspondences is not well defined outside of the extendible setting.